# D-76131 Karlsruhe, Germany

### v1

@@ 1 Introduction
The phenomenon of neutrino oscillations is, by now, a textbook subject, see e.g. refs. giunti ; Suekane:2015yta . Nevertheless, discussions of its proper treatment and interpretation both in quantum mechanics and, especially, in quantum field theory (QFT) continue, see for instance refs. Kayser:1981ye ; Giunti:1991ca ; Giunti:1993se ; Grimus:1996av ; Kiers:1997pe ; Beuthe:2001rc ; Cohen:2008qb ; Naumov:2010um ; Grimus:2019hlq ; Naumov:2020yyv . This reflects an obvious fact that neutrino oscillations is a macroscopic phenomenon, whereas the apparatus of quantum field theory was developed to address physics of microscopic origin, and many familiar concepts in QFT, such as cross sections and decay rates, are designed to serve this purpose. In principle, from the point of view of quantum field theory, the description of neutrino oscillations should be straightforward. One considers a single scattering amplitude where a neutrino is first produced in a particular weak decay process, then propagates and eventually collides with a “detector particle” producing a final state which allows its flavour identification.
However, relating the scattering amplitude and the neutrino oscillation probability must involve a modification of conventional computational methods. Indeed, this probability depends on the distance between the source of neutrinos and the detector, and distances do not appear in calculations of standard cross sections in QFT. This conundrum was resolved thirty years ago, when it was recognized Rich:1993wu ; Kiers:1995zj ; Grimus:1996av ; Giunti:1997sk ; Giunti:1997wq that one should employ wave packets for external particles to describe neutrino oscillations in QFT. Unfortunately, since working with wave packets is significantly more involved than with plane waves, one often attempts to simplify calculations of the quantum amplitude by assuming a particular (usually Gaussian) form of wave packets. Although this approach leads to important insights into the physics of neutrino oscillations, it may hide the generality of the obtained results Grimus:1996av ; Beuthe:2001rc ; Giunti:1993se ; Grimus:2019hlq ; Kiers:1997pe ; Cohen:2008qb ; Naumov:2010um ; Naumov:2020yyv .
@@ 2 Derivation of the neutrino oscillation probability
It is well-appreciated by now Grimus:1996av ; Beuthe:2001rc ; Giunti:1993se ; Grimus:2019hlq ; Kiers:1997pe ; Naumov:2010um ; Naumov:2020yyv that establishing a connection between the scattering amplitude and the neutrino oscillation probability, requires introduction of wave packets for particles $S$ and $D$ in the initial state. Quite often wave packets for the final-state particles are also introduced but, as we will see, this is not necessary and we will not do that in what follows.
@@ 5 Summary and discussion
Our approach enables the derivation of the standard oscillation formula without the need to specify the shape of the initial-state wave packets. Instead of working with transition amplitudes, we consider the transition probability (proportional to the amplitude-squared) from the very beginning. A statistical average of the probability leads to density matrices for neutrino source and detector particles, which in turn are expressed in terms of semi-classical phase-space densities with the help of the Wigner transformation Case2008 . This approach offers a very transparent way to take into account properties of neutrino source and detector. One of our main results is eq. 40, where the transition rate is obtained as a convolution of the production rate, the oscillation probability and the detection cross section with the source and detector phase-space densities.
@@
A crucial assumption of our method is that all final-state particles are exact momentum eigenstates. The model of ref. Beuthe:2001rc used e.g., in ref. Krueger:2023skk would lead to a vanishing interference term, in the case of a single source and detector particle and sharp final-state momenta. We believe that this is a consequence of expanding the dispersion relation only up to linear order combined with the Gaussian wave-packet ansatz, which leads to an intricate cancellation of terms setting the effective energy spread to zero. This is different from our model, which allows for interference under these assumptions. We have checked indeed that kinematics does not fully determine the neutrino momentum under these conditions, and therefore, an interference term is not in contradiction with the energy-momentum conservation implied by the sharp final-state momenta.
The last item may have also phenomenological consequences. Contrary to our assumption that momenta of final-state particles are fixed, in the approach of refs. Beuthe:2001rc ; Akhmedov:2010ms ; Krueger:2023skk their uncertainties are taken into account. It is found that these uncertainties impact the coherence properties. In fact, it turns out, that the effective energy-smearing is often dominated by the outgoing charged lepton, which typically has a large mean-free path and therefore a fairly well-defined momentum. This leads to values of $\sigma$ which can be ${\cal O}(10^{3})$ times smaller than what one obtains by considering initial-state particles, see e.g. table 1 in ref. Krueger:2023skk , or ref. Akhmedov:2022bjs for similar results based on the wave-packet arguments. Our assumption of fixed final-state momenta is rooted in the standard particle physics calculation of scattering rates and cross sections. We leave the question of whether for certain experimental configurations this approach needs to be adapted to correctly describe phenomenology of neutrino oscillations for further studies.

### v2

@@ 1 Introduction
The phenomenon of neutrino oscillations is, by now, a textbook subject, see e.g. refs. giunti ; Suekane:2015yta . Nevertheless, discussions of its proper treatment and interpretation both in quantum mechanics and, especially, in quantum field theory (QFT) continue, see for instance refs. Kayser:1981ye ; Giunti:1991ca ; Giunti:1993se ; Grimus:1996av ; Kiers:1997pe ; Ioannisian:1998ch ; Beuthe:2001rc ; Cohen:2008qb ; Falkowski:2019kfn ; Grimus:2019hlq ; Naumov:2020yyv ; Karamitros:2022nnh . This reflects an obvious fact that neutrino oscillations is a macroscopic phenomenon, whereas the apparatus of quantum field theory was developed to address physics of microscopic origin, and many familiar concepts in QFT, such as cross sections and decay rates, are designed to serve this purpose. In principle, from the point of view of quantum field theory, the description of neutrino oscillations should be straightforward. One considers a single scattering amplitude where a neutrino is first produced in a particular weak decay process, then propagates and eventually collides with a “detector particle” producing a final state which allows its flavour identification.
However, relating the scattering amplitude and the neutrino oscillation probability must involve a modification of conventional computational methods. Indeed, this probability depends on the distance between the source of neutrinos and the detector, and distances do not appear in calculations of standard cross sections in QFT. This conundrum was resolved thirty years ago, when it was recognized Giunti:1993se ; Kiers:1995zj ; Grimus:1996av ; Giunti:1997sk ; Giunti:1997wq that one should employ wave packets for external particles to describe neutrino oscillations in QFT. Unfortunately, since working with wave packets is significantly more involved than with plane waves, one often attempts to simplify calculations of the quantum amplitude by assuming a particular (usually Gaussian) form of wave packets. Although this approach leads to important insights into the physics of neutrino oscillations, it may hide the generality of the obtained results Grimus:1996av ; Ioannisian:1998ch ; Giunti:1993se ; Grimus:2019hlq ; Kiers:1997pe ; Cohen:2008qb ; Naumov:2010um ; Naumov:2020yyv ; Karamitros:2022nnh .
@@ 2 Derivation of the neutrino oscillation probability
It is well-appreciated by now Grimus:1996av ; Beuthe:2001rc ; Giunti:1993se ; Grimus:2019hlq ; Kiers:1997pe ; Naumov:2010um ; Naumov:2020yyv that establishing a connection between the scattering amplitude and the neutrino oscillation probability, requires introduction of wave packets for particles $S$ and $D$ in the initial state. Quite often wave packets for the final-state particles are also introduced but, as we will see, this is not necessary and we will not do that in what follows, see also Grimus:1996av ; Falkowski:2019kfn ; Grimus:2019hlq .
@@ 5 Summary and discussion
Our approach enables the derivation of the standard oscillation formula without the need to specify the shape of the initial-state wave packets. Instead of working with transition amplitudes, we consider the transition probability (proportional to the amplitude-squared) from the very beginning. These aspects of our derivation are similar to the one presented in appendix A of Falkowski:2019kfn , arXiv version v3. There, the interference term is obtained from a pole-integration, which has similarities to our $\mathrm{d}^{4}\kappa$ integration, see eq. 27. The latter generalizes the corresponding expression in Falkowski:2019kfn to four space-time dimensions, allowing for a consistent discussion of time dependencies, as we discuss below.
Following refs. Ginzburg:1991mv ; Ginzburg:1991mw ; Kotkin:1992bj , we consider a statistical average of the probability which leads to density matrices for neutrino source and detector particles, which in turn are expressed in terms of semi-classical phase-space densities with the help of the Wigner transformation Case2008 . This approach offers a very transparent way to take into account properties of neutrino source and detector. One of our main results is eq. 40, where the transition rate is obtained as a convolution of the production rate, the oscillation probability and the detection cross section with the source and detector phase-space densities.
@@
A crucial assumption of our method is that all final-state particles are exact momentum eigenstates, see also Grimus:1996av ; Falkowski:2019kfn ; Grimus:2019hlq . The model of ref. Beuthe:2001rc used e.g., in ref. Krueger:2023skk would lead to a vanishing interference term, in the case of a single source and detector particle and sharp final-state momenta. We believe that this is a consequence of expanding the dispersion relation only up to linear order combined with the Gaussian wave-packet ansatz, which leads to an intricate cancellation of terms setting the effective energy spread to zero. This is different from our model, which allows for interference under these assumptions. We have checked indeed that kinematics does not fully determine the neutrino momentum under these conditions, and therefore, an interference term is not in contradiction with the energy-momentum conservation implied by the sharp final-state momenta.
The last item may have also phenomenological consequences. Contrary to our assumption that momenta of final-state particles are fixed, in the approach of refs. Giunti:1993se ; Beuthe:2001rc ; Akhmedov:2010ms ; Krueger:2023skk their uncertainties are taken into account. It is found that these uncertainties impact the coherence properties. In fact, it turns out, that the effective energy-smearing is often dominated by the outgoing charged lepton, which typically has a large mean-free path and therefore a fairly well-defined momentum. This leads to values of $\sigma$ which can be ${\cal O}(10^{3})$ times smaller than what one obtains by considering initial-state particles, see e.g. table 1 in ref. Krueger:2023skk , or ref. Akhmedov:2022bjs for similar results based on the wave-packet arguments. Our assumption of fixed final-state momenta is rooted in the standard particle physics calculation of scattering rates and cross sections. We leave the question of whether for certain experimental configurations this approach needs to be adapted to correctly describe phenomenology of neutrino oscillations for further studies.